A 1290-kg car pulls a 270-kg trailer. The car exerts a horizontal force of 3.6 × 103 N against the ground in order to accelerate.
What force does the car exert on the trailer? Assume an effective friction coefficient of 0.16 for the trailer.
1
Expert's answer
2019-02-19T14:53:08-0500

The Newton’s second law for a car

Ma=F-T

The Newton’s second law for a trailer

ma=T-F_frict=T-μmg

By combining two equations, we obtain

T=(m(F+μMg))/(m+M)

=(270(3.6×〖10〗^3+0.16×1290×9.81))/(270+1290)=974 N